how many seconds does it take to travel 500 feet at 30 miles per hour?
Answer by mananth(16946)

30 mph
=30*(5280/3600) feet per second
1 miles 5280 feet
1 hour = 3600 seconds
= 44 feet /second
t=d/r
d=500
r= 44 ft/s
t= 500/44
11.36 seconds